Steve can paint a room in 3.5 hours, while working together with Jason they will paint the room in 2 hours. How much time would it take Jason to paint the room alone?

Answer by checkley79(3341)

1/3.5+1/X=1/2
1/X=1/2-1/3.5
1/X=(3.5-2)/2*3.5
1/X=1.5/7 CROSS MULTIPLY
1.5X=7
X=7/1.5
X=4 2/3 HOURS FOR JASON WORKING ALONE.
PROOF:
1/3.5+1/(4 2/3)=1/2
1/3.5+1/(12+2)/3=1/2
1/3.5+(1*3/14)=1/2
1/3.5+3/14=1/2
(4+3)/14=1/2
1/2=1/2